5. The nth term of an A.P. 5, 2, -1, -4, -7 … is
(a) 2n + 5
(b) 2n – 5
(c) 8 – 3n
(d) 3n – 8

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: c Explaination:Reason: Here a = 5, d = 2 – 5 = -3 an = a + (n – 1)d = 5 + (n – 1) (-3) = 5 – 3n + 3 = 8 – 3n